# Integrated Radiation-Magneto-Hydrodynamic Simulations of Magnetized Burning Plasmas. I. Magnetizing Ignition-Class Designs

### v1

@@ IV Basic Physics of Magnetized Hohlraums
The application of 10-100 T magnetic fields onto an ICF target is a non-trivial physical and engineering problem. For example, magnetized ICF experiments on NIF necessitated the development of a new gold-tantalum alloy hohlraum which has comparable X-ray generation as pure gold but an increased electrical resistivity so that it is not crushed by the applied field.moody1 For a magnetic field to have a perturbative effect on an ICF hotspot plasma, kT-scale fields are necessary. However, this is actually a relatively simple goal to achieve given the properties of an ICF implosion. An ICF capsule, when ideally compressed, will have a convergence ratio of 30, i.e., initial/final capsule radius, which means that the area of a magnetic flux surface that spans the diameter of the capsule will be compressed by a factor of 900. Therefore, an applied 30 T field, which has been almost achieved in the NIF subscale, symcap experiments thus far,sio1; moody1; strozzi1 will ideally reach 27 kT scales at stagnation. This field is large enough to have an impact even with some flux loss, e.g., through resistive diffusion and the Nernst effect.
Flux compression can be described by Alfvén’s theorem, which says that magnetic flux in a conducting fluid is conserved through deformations of material surfaces, i.e., , where is the magnetic flux and is the advective or material derivative.blackman In an ideal plasma that undergoes a reduction in radius from to and where magnetic flux remains constant, , where is the surface area in question, the magnetic field scales as:
This process is governed by the induction equation (here simplified versus to the one implemented in Lasnex):
@@
where is the unit vector along B, and is the Hall (Righi-Leduc) conductivity giving a heat flux perpendicular to both and . The thermal conductivities in eqn. 4 scale as follows:braginskii1965
@@
The second benefit of magnetic fields is that they are believed to reduce Rayleigh-Taylor instabilities. The basic concept here is that field lines resist perturbations perpendicular to them, i.e., there is a restoring force against the bending of field lines, . This has yet to be demonstrated in integrated ICF experiments but has been seen in simulations.perkins2; walsh2 The idea that hydrodynamic mix can be reduced via magnetization comes from the concept of magnetic tension, which acts to straighten out bent field lines. The actual dynamics of the evolution of the magnetic field inside the capsule is complicated by various MHD effects, such as the Nernst effect that transports perpendicular field lines down temperature gradients per equations 2-3, i.e., on the equator. However, the primary perturbation to spherical field compression is that of hydrodynamic ablation, which pinches in field lines further inward on the equator as opposed to the poles as the ice layer heats up and ablates into the hotspot.
@@
The primary object of this study is that of the implosion itself. However, in order to implode a capsule we need a hohlraum, and we need to validate that we may treat magnetization of the capsule as being relatively independent of the hohlraum. A recent study of fully-integrated Lasnex simulations modeling gas-filled symcaps suggested that magnetization at T initially does not significantly affect the physics of the hohlraum,strozzi1 namely the radiation temperature generated by the laser-heated walls of the hohlraum (there is a modest effect on x-ray flux asymmetry per Lasnex modeling). This has also been observed in experiments, where the observed when magnetized at T is almost identical to when unmagnetized.sio1 We verify this for all igniting designs by looking at as a function of time, which is an effective metric of the x-ray drive experienced by the capsule. is practically identical for the duration of the entire implosion in Fig. 4 (a)-(e) for all five designs. The variation after bangtime is typical even when unmagnetized and thereby not a concern. The only exception is for PSS and EYC, where there is a slight decrease in late in time, particularly for EYC, but this is something that can be corrected for in future designs.
@@
In addition, we can extract a Legendre decomposition of the x-ray flux in the hohlraum to get an idea of the degree of asymmetry present in the simulation, as seen in Fig. 4 (f) - (o). The divergence of when magnetized seems small relative to when unmagnetized and differences are neglible, but shape control is challenging and variable on its own even when unmagnetized. An applied magnetic field may alter CBET both directly, by changing the basic CBET coupling (currently not included in Lasnex or any rad-MHD code we are familiar with), and indirectly, by changing the LEH plasma conditions (included in Lasnex, to the extent the MHD model is valid). For magnetized simulations, we use the CBET clamp tuned to match the relevant unmagnetized data. Even if this is not accurate, small changes in the imposed wavelength shift should be able to tune magnetized implosion shape similarly to unmagnetized cases. In addition, the evolution of the gold bubble on the hohlraum wall is relatively unaffected. Earlier work showed measurements of the electron temperature in the gold bubble of NIF hohlraums only match Lasnex modeling when the heat flow is restricted compared to the unmagnetized Spizter-Härm level.meezan2020 The Lasnex MHD model naturally provides this, if the Nernst effect is multiplied by .rosen2018 We follow this and use outside the capsule, but 1 inside it. More recent modeling work with modified gold properties (namely opacity and heat capacity) allows us to keep in the hohlraum to agree better with NIF data.chen2024
@@ V.1 N180128 - highest performing BigFoot shot
As discussed in the introduction, a technique we use is symmetrization, whereby we can mimic the ideal, 1D result in 2D. This can be seen in Fig. 5(a) with the dashed blue and red lines for symmetrized yield and , respectively. In this case the maximum magnetic enhancment of 2.1 can be found at 50 T instead of 30 T. In addition, when symmetrized, the relative enhancement due to magnetization rises much more rapidly for small, applied magnetic fields. Contrary to expectations, can actually be lower generally for applied magnetic fields when symmetrized. It has been suggested that for clearly non-igniting implosions, magnetic pressure can be a more prominent effect,ho2 preventing the implosion from compressing as much, particularly on the equator.
@@ V.2 N210808 - Lawson criterion shot
A significant advance in the LLNL ICF program occurred in 2021, when NIF shot N210808 exceeded a fusion yield of 1 MJ for the first time.abu1 This shot was part of the HYBRID-E campaign,hurricane which built upon BigFoot and other campaigns with HDC ablators. Its goal was to increase yield by coupling more energy to a larger capsule and pushing hohlraum performance further towards its limit, e.g., a smaller LEH.
@@
Performing a magnetic field strength scan, we see that the yield increases significantly, by a factor >6 for T, as seen in Fig. 5 (b). However, the burn-averaged does not increase as dramatically, with just a 50% increase in line with what was previously seen for N180128. N210808 is distinct in that the symmetrized, magnetized yield is lower than the unsymmetrized, magnetized yield for lower applied field strengths (but still higher than the unmagnetized yield) except near T, where the enhancement goes up from 6 to 8. This may suggest that the asymmetric nature of the implosion, such as the polar caps, was actually more conducive to yield enhancement. However, the compression of a symmetrized implosion is further enhanced that at the point of peak magnetic enhancement.
@@
To get a better understanding of the implosion we can again look at lineouts at bangtime (9.26 ns), as seen in Fig. 6 (b), comparing T and T. Note the lineouts in Fig. 6 (b) are normalized differently than in Fig. 6 (a). This time the profiles look notably different. The implosion is actually less converged when magnetized on the equator but on the poles the density profile has been completely blown through in a highly asymmetric fashion. One can see that not only is the temperature almost 3 higher at this time but that there is significant temperature disequilibration between the ions and electrons. On the pole an expansion shock has formed, which is characterized by high ion temperatures relative to that of the electrons. This is due to hydrodynamic viscosity effects independent of magnetization,velikovich where shocks preferentially heat ions. In a robustly igniting target, this occurs not only during the implosion but also during the detonation phase, although there are additional second-order effects of magnetization on shock dynamics as well. Again, magnetic field strengths approach 50 kT particularly near the gas-ice interface, where internal ablation of fuel layer piles up the B-field inward.
@@
To highlight the evolution of the magnetized case example we have plotted the density, temperature, and magnetic field profiles at sequential times between peak compression and bangtime, in Fig. 9 (a) for T and Fig. 9 (b) for T where magnetic enhancement is greatest. We can see the clear formation of polar caps in both cases, typical of current NIF designs, but in the magnetized case one can clearly see the polar cap blowing out early relative to bangtime at 9.26 ns, which seems to be common occurrence in high yield, magnetized designs. This is reasonable given that there are greater heat conduction and alpha propagation on the poles versus the equator. The unusual polar burn-through of hotspot profile at bangtime can be readily explained by examining the implosion several tens of picoseconds before bangtime, i.e., at 9.23 ns when unmagnetized and 9.22 ns when magnetized. In Fig. 9 (a) and (b) we can see 2D images of the density for T and T at relative times and we can observe that the capsule shell is still intact for the magnetized case. In both cases we see the denser polar caps that are typical of HYBRID-E designs born out of the and swings inherent to laser-driven hohlraums. These swings are due to the LEH closure affecting laser propagation and the gold bubble driven by the outer laser cones absorbing energy from the inner cones. Symmetrized results can be seen in Fig. 9 (c), where we can see the presence of magnetic fields result in a -like structure in the density and temperature profiles.
@@ V.3 N221204 - first breakeven shot
In 2022, NIF shot N221204 achieved energy breakeven for the first time,abu2 meaning that more energy was produced via nuclear reactions than delivered by the drive laser to the hohlraum. The most notable change from N210808 was that the laser energy was increased from 1.9 MJ to 2.05 MJ, slightly lengthening the peak laser drive with respect to time. In addition, the thickness of the capsule ablator was slightly increased, beam pointing and were retuned to control implosion shape. As noted before, our prediction of the yield was somewhat high, 4.6 MJ versus the experimental 3.2 MJ, but we kept this result since recent experiments based on the N221204 design have since exceeded the original yield studied here.
@@
With symmetrization N221204 doubles in yield even without an applied magnetic field, suggesting that shape improvements or reduction in residual kinetic energy could significantly enhance the yield. Mix, however, was a minor factor, given that the no-mix, symmetrized result does not visibly differ. However, the peak enhancement still comes with magnetization, reaching in yield enhancement at 70 T with respect to the unmagnetized, unsymmetrized data point but only a 50% increase relative to the unmagnetized, symmetrized zero point. This suggests that laser drive symmetry can be improved even further without an applied magnetic field but that it is still beneficial to apply one. However, achieving such a perfect implosion is currently impossible, as the and swings in the x-ray drive will result in some asymmetry, i.e., polar caps. It is possible that magnetic fields can synergies with such asymmetries however.
@@ VI Pushered Single Shell (PSS)
PSS is an alternative approach to HYBRID-E whereby high-Z materials are used in the inner capsule layer in order to enhance confinement and radiation trapping and lower ignition requirements. In PSS we have a relatively dense, thick shell that has a density gradient throughout its radius as previously depicted in Fig. 1. The shape of the Mo concentration is given by,maclaren3
@@
We first perform a magnetic field scan for PSS with a 1.9 MJ laser drive and medium mix. We see a pronounced 4 peak in yield enhancement at 50 T as seen in Fig. 5 when unsymmetrized and 12.5 at 30 T when symmetrized, going from 63 kJ to 253 kJ and 788 kJ, respectively. For both moderate and low mix, a visualization of the time evolution can be seen in Figs. 14 and 15. When lowering mix levels by 20 %, we see significant enhancement in yield both when magnetized and unmagnetized. For an asymmetric drive we see a peak magnetic enhancement at 30 T while for symmetrized drives we have a peak at 30 T, going from 0.97 MJ to 6.3 MJ (6.5) and to 9.7 MJ (10) for asymmetric and symmetric drives, respectively. The significant enhancement due to symmetrization suggests that a PSS design is a promising candidate for further optimization, although in addition to mix it was also found to be more sensitive to shape asymmetries than HYBRID-E designs. However, these are baseline design questions that do not preclude future magnetization of them, and there exists the possibility that asymmetries caused by magnetization may balance out those caused by hydrodynamic asymmetries.
@@ VII Enhanced Yield Capability
Ongoing efforts in the ICF community have been looking at enhancing the energy of the NIF laser from 2.2 MJ to 2.6 MJ by the 2030s, with potential for going up to 3 MJ thereafter. In this study we consider a 3 MJ EYC design as presented in Ref. maclaren1. In this design, the hohlraum is wider and longer given the longer laser and has a larger LEH than current high-performing designs. This is necessary to control symmetry for the larger capsule as well as prevent Au LEH plasma blow-off from interfering with the 44∘ and 50∘ laser beams. There is a higher picket at the beginning of the laser pulse which complicates symmetry. The injection of high power at the beginning of the laser pulse is needed for the first shock to completely melt the HDC ablator. However, this causes faster growth of the gold bubble on the hohlraum wall, causing more of the inner cone power to be absorbed closer to the outer beam spots, pushing the capsule symmetry more towards an oblate implosion.callahan1 This is counteracted by applying greater wavelength separation between the inner (23∘,30∘) and outer (44∘,50∘) beams, going from Å to Å. A challenge with the 3 MJ HDC design used in this study is that one must overcome radiative losses caused by mixing of the carbon ablator into the hotspot. This requires a higher implosion velocity than necessary for previous designs. Increased velocity provides the necessary hotspot energy and density to overcome bremsstrahlung losses. A peak fuel velocity close to that of N221204 was achieved by increasing the increasing the ablator thickness and reducing the fuel mass.
Magnetizing the 3 MJ EYC design demonstrates the potential limitations of simple application of external magnetic fields. As seen in Fig. 5 (f), the benefit of magnetization for this design is limited if not detrimental. First of all, we were not able to retrieve the full 25 MJ results from Ref. maclaren1, achieving at best an initial 13.7 MJ yield at T, which steadily degrades with increased magnetic field strength. The primary difference here is that we used power multipliers as opposed to opacity multipliers to be consistent with the rest of the simulations in this study. When symmetrized, which is the ideal result, we are able to achieve 30 MJ. We can see the negligible influence of a magnetic field up to 100 T, with slight degradation in the yield until it rises again at 90 T, before falling precipitously again as we increase field strength up to 200 T. The peak in at 90 T suggests that we might obtain some modest benefit of magnetization with some optimization of the design.
@@
As yield performance increases, we see diminishing returns from applied magnetic fields. The basic understanding is that the hotspot already gets so hot that it does not heat much more from constrained heat conduction and alpha magnetization. There is a natural peak in the cross section of the DT nuclear reaction at about 50 keV, and we see clearly diminishing returns as one exceeds 20 keV. This trend can be seen in Fig. 17, where we plot the burn fraction, i.e., , versus an analytical modelatzeni of burn efficiency.
for the shots in consideration (empty circle), the peak-associated magnetization case (empty star), and the associated symmetrized runs (solid circle and star). Here and is the mass and areal density of the imploded fuel, respectively, as it begins to burn and . In addition, we have plotted out the burn-efficiency curves for constant values, while for the shots in question we considered the maximum achieved in the simulation. The estimated maximum burn efficiency is 30%, which is being approached by N221204, PSS with low mix, and EYC when magnetized. When designs are already achieving very hot temperatures they will not benefit from magnetization, thus the motivation for the PSS design.

### v2

@@ IV Basic Physics of Magnetized Hohlraums
The application of 10-100 T magnetic fields onto an ICF target is a non-trivial physical and engineering problem. For example, magnetized ICF experiments on NIF necessitated the development of a new gold-tantalum alloy hohlraum which has comparable X-ray generation as pure gold but an increased electrical resistivity so that it is not crushed by the applied field.Moody et al. (2022) For a magnetic field to have a perturbative effect on an ICF hotspot plasma, kT-scale fields are necessary. However, this is actually a relatively simple goal to achieve given the properties of an ICF implosion. An ICF capsule, when ideally compressed, will have a convergence ratio of 30, i.e., initial/final capsule radius, which means that the area of a magnetic flux surface that spans the diameter of the capsule will be compressed by a factor of 900. Therefore, an applied 30 T field, which has been almost achieved in the NIF subscale, symcap experiments thus far,Sio et al. (2023); Moody et al. (2022); Strozzi et al. (2024) will ideally reach 27 kT scales at stagnation. This field is large enough to have an impact even with some flux loss, e.g., through resistive diffusion and the Nernst effect.
Flux compression can be described by Alfvén’s theorem, which says that magnetic flux in a conducting fluid is conserved through deformations of material surfaces, i.e., , where is the magnetic flux and is the advective or material derivative.Blackman (2013) In an ideal plasma that undergoes a reduction in radius from to and where magnetic flux remains constant, , where is the surface area in question, the magnetic field scales as:
This process is governed by the induction equation (here simplified versus the one implemented in Lasnex):
@@
where is the unit vector along B, and is the Hall (Righi-Leduc) conductivity giving a heat flux perpendicular to both and . The thermal conductivities in Eqn. 4 scale as follows:Braginskii (1965)
@@
The second benefit of magnetic fields is that they are believed to reduce Rayleigh-Taylor instabilities. The basic concept here is that field lines resist perturbations perpendicular to them, i.e., there is a restoring force against the bending of field lines, . This has yet to be demonstrated in integrated ICF experiments but has been seen in simulations.Perkins et al. (2017); Walsh et al. (2024) The idea that hydrodynamic mix can be reduced via magnetization comes from the concept of magnetic tension, which acts to straighten out bent field lines. The actual dynamics of the evolution of the magnetic field inside the capsule is complicated by various MHD effects, such as the Nernst effect that transports perpendicular field lines down temperature gradients per equations 2-3, i.e., on the equator. However, the primary perturbation to spherical field compression is that of hydrodynamic ablation, which pinches in field lines further inward on the equator as opposed to the poles as the ice layer heats up and ablates into the hotspot.
@@
The primary object of this study is that of the implosion itself. However, in order to implode a capsule we need a hohlraum, and we need to validate that we may treat magnetization of the capsule as being relatively independent of the hohlraum. A recent study of fully-integrated Lasnex simulations modeling gas-filled symcaps suggested that magnetization at T initially does not significantly affect the physics of the hohlraum,Strozzi et al. (2024) namely the radiation temperature generated by the laser-heated walls of the hohlraum (there is a modest effect on x-ray flux asymmetry per Lasnex modeling). This has also been observed in experiments, where the observed when magnetized at T is almost identical to when unmagnetized.Sio et al. (2023) We verify this for all igniting designs by looking at as a function of time, which is an effective metric of the x-ray drive experienced by the capsule. is practically identical for the duration of the entire implosion in Fig. 4 (a)-(e) for all five designs. The variation after bangtime is typical even when unmagnetized and thereby not a concern. The only exception is for PSS and EYC, where there is a slight decrease in late in time, particularly for EYC, but this is something that can be corrected for in future designs.
@@
In addition, we can extract a Legendre decomposition of the x-ray flux in the hohlraum to get an idea of the degree of asymmetry present in the simulation, as seen in Fig. 4 (f) - (o). The divergence of when magnetized seems small relative to when unmagnetized and differences are neglible, but shape control is challenging and variable on its own even when unmagnetized. An applied magnetic field may alter CBET both directly, by changing the basic CBET coupling (currently not included in Lasnex or any rad-MHD code we are familiar with), and indirectly, by changing the LEH plasma conditions (included in Lasnex, to the extent the MHD model is valid). For magnetized simulations, we use the CBET clamp tuned to match the relevant unmagnetized data. Even if this is not accurate, small changes in the imposed wavelength shift should be able to tune magnetized implosion shape similarly to unmagnetized cases. In addition, the evolution of the gold bubble on the hohlraum wall is relatively unaffected. Earlier work showed measurements of the electron temperature in the gold bubble of NIF hohlraums only match Lasnex modeling when the heat flow is restricted compared to the unmagnetized Spizter-Härm level.Meezan et al. (2020) The Lasnex MHD model naturally provides this, if the Nernst effect is multiplied by .Rosen (2018) We follow this and use outside the capsule, but 1 inside it. More recent modeling work with modified gold properties (namely opacity and heat capacity) allows us to keep in the hohlraum to agree better with NIF data.Chen et al. (2024)
@@ V.1 N180128 - highest performing BigFoot shot
As discussed in the introduction, a technique we use is symmetrization, whereby we can mimic the ideal, 1D result in 2D. This can be seen in Fig. 5(a) with the dashed blue and red lines for symmetrized yield and , respectively. In this case the maximum magnetic enhancment of 2.1 can be found at 50 T instead of 30 T. In addition, when symmetrized, the relative enhancement due to magnetization rises much more rapidly for small, applied magnetic fields. Contrary to expectations, can actually be lower generally for applied magnetic fields when symmetrized. It has been suggested that for clearly non-igniting implosions, magnetic pressure can be a more prominent effect,Zimmerman et al. (2014) preventing the implosion from compressing as much, particularly on the equator.
@@ V.2 N210808 - Lawson criterion shot
A significant advance in the LLNL ICF program occurred in 2021, when NIF shot N210808 exceeded a fusion yield of 1 MJ for the first time.Abu-Shawareb et al. (2022) This shot was part of the HYBRID-E campaign,Hurricane et al. (2018) which built upon BigFoot and other campaigns with HDC ablators. Its goal was to increase yield by coupling more energy to a larger capsule and pushing hohlraum performance further towards its limit, e.g., a smaller LEH.
@@
Performing a magnetic field strength scan, we see that the yield increases significantly, by a factor >6 for T, as seen in Fig. 5 (b). However, the burn-averaged does not increase as dramatically, with just a 50% increase in line with what was previously seen for N180128. N210808 is distinct in that the symmetrized, magnetized yield is lower than the unsymmetrized, magnetized yield for lower applied field strengths (but still higher than the unmagnetized yield) except near T, where the enhancement goes up from 6 to 8. This may suggest that the asymmetric nature of the implosion, such as the polar caps, was actually more conducive to yield enhancement. However, the compression of a symmetrized implosion is still the superior result when considering peak magnetic enhancement.
@@
To get a better understanding of the implosion we can again look at lineouts at bangtime (9.26 ns), as seen in Fig. 6 (b), comparing T and T. Note the lineouts in Fig. 6 (b) are normalized differently than in Fig. 6 (a). This time the profiles look notably different. The implosion is actually less converged when magnetized on the equator but on the poles the density profile has been completely blown through in a highly asymmetric fashion. One can see that not only is the temperature almost 3 higher at this time but that there is significant temperature disequilibration between the ions and electrons. On the pole an expansion shock has formed, which is characterized by high ion temperatures relative to that of the electrons. This is due to hydrodynamic viscosity effects independent of magnetization,Velikovich et al. (2001) where shocks preferentially heat ions. In a robustly igniting target, this occurs not only during the implosion but also during the detonation phase, although there are additional second-order effects of magnetization on shock dynamics as well. Again, magnetic field strengths approach 50 kT particularly near the gas-ice interface, where internal ablation of the fuel layer piles up the B-field inward.
@@
To highlight the evolution of the magnetized case example we have plotted the density, temperature, and magnetic field profiles at sequential times between peak compression and bangtime, in Fig. 9 (a) for T and Fig. 9 (b) for T where magnetic enhancement is greatest. We can see the clear formation of polar caps in both cases, typical of current NIF designs, but in the magnetized case one can clearly see the polar cap blowing out early relative to bangtime at 9.26 ns, which seems to be a common occurrence in high yield, magnetized designs. This is reasonable given that there are greater heat conduction and alpha propagation on the poles versus the equator. The unusual polar burn-through of hotspot profile at bangtime can be readily explained by examining the implosion several tens of picoseconds before bangtime, i.e., at 9.23 ns when unmagnetized and 9.22 ns when magnetized. In Fig. 9 (a) and (b) we can see 2D images of the density for T and T at relative times and we can observe that the capsule shell is still intact for the magnetized case. In both cases we see the denser polar caps that are typical of HYBRID-E designs born out of the and swings inherent to laser-driven hohlraums. These swings are due to the LEH closure affecting laser propagation and the gold bubble driven by the outer laser cones absorbing energy from the inner cones. Symmetrized results can be seen in Fig. 9 (c), which show that magnetic fields result in a -like structure in the density and temperature profiles.
@@ V.3 N221204 - first breakeven shot
In 2022, NIF shot N221204 achieved energy breakeven for the first time,Abu-Shawareb et al. (2024) meaning that more energy was produced via nuclear reactions than delivered by the drive laser to the hohlraum. The most notable change from N210808 was that the laser energy was increased from 1.9 MJ to 2.05 MJ, slightly lengthening the peak laser drive with respect to time. In addition, the thickness of the capsule ablator was slightly increased, beam pointing and were retuned to control implosion shape. As noted before, our prediction of the yield was somewhat high, 4.6 MJ versus the experimental 3.2 MJ, but we kept this result since recent experiments based on the N221204 design have since exceeded the original yield studied here.
@@
With symmetrization N221204 doubles in yield even without an applied magnetic field, suggesting that shape improvements or reduction in residual kinetic energy could significantly enhance the yield. Mix, however, was a minor factor, given that the no-mix, symmetrized result does not visibly differ. However, the peak enhancement still comes with magnetization, reaching in yield enhancement at 70 T with respect to the unmagnetized, unsymmetrized data point but only a 50% increase relative to the unmagnetized, symmetrized zero point. This suggests that laser drive symmetry can be improved even further without an applied magnetic field but that it is still beneficial to apply one. However, achieving such a perfect implosion is currently impossible, as the and swings in the x-ray drive will result in some asymmetry, i.e., polar caps. It is possible that magnetic fields can synergize with such asymmetries however.
@@ VI Pushered Single Shell (PSS)
PSS is an alternative approach to HYBRID-E whereby high-Z materials are used in the inner capsule layer in order to enhance confinement and radiation trapping and lower ignition requirements. In PSS we have a relatively dense, thick shell that has a density gradient throughout its radius as previously depicted in Fig. 1. The shape of the Mo concentration is given by,MacLaren et al. (2025)
@@
We first perform a magnetic field scan for PSS with a 1.9 MJ laser drive and medium mix, which is closer to what has been observed in radiochemistry experiments thus far. We see a pronounced 4 peak in yield enhancement at 50 T as seen in Fig. 5 when unsymmetrized and 12.5 at 30 T when symmetrized, going from 63 kJ to 253 kJ and 788 kJ, respectively. For both moderate and low mix, a visualization of the time evolution can be seen in Figs. 14 and 15. When lowering mix model parameters by 20 %, we see significant enhancement in yield both when magnetized and unmagnetized. For an asymmetric and symmetric drives we see a peak magnetic enhancement at 30 T, going from 0.97 MJ to 6.3 MJ (6.5) and to 9.7 MJ (10) for asymmetric and symmetric drives, respectively. The significant enhancement due to symmetrization suggests that a PSS design is a promising candidate for further optimization, although in addition to mix it was also found to be more sensitive to shape asymmetries than HYBRID-E designs. However, these are baseline design questions that do not preclude future magnetization of them, and there exists the possibility that asymmetries caused by magnetization may balance out those caused by hydrodynamic asymmetries.
@@ VII Enhanced Yield Capability
Ongoing efforts in the ICF community have been looking at enhancing the energy of the NIF laser from 2.2 MJ to 2.6 MJ by the 2030s, with potential for going up to 3 MJ thereafter. In this study we consider a 3 MJ EYC design as presented in Ref. MacLaren et al., 2024. In this design, the hohlraum is wider and longer given the longer laser pulse and has a larger LEH than current high-performing designs. This is necessary to control symmetry for the larger capsule as well as prevent Au LEH plasma blow-off from interfering with the 44∘ and 50∘ laser beams. There is a higher picket at the beginning of the laser pulse which complicates symmetry. The injection of high power at the beginning of the laser pulse is needed for the first shock to completely melt the HDC ablator. However, this causes faster growth of the gold bubble on the hohlraum wall, causing more of the inner cone power to be absorbed closer to the outer beam spots, pushing the capsule symmetry more towards an oblate implosion.Callahan et al. (2018) This is counteracted by applying greater wavelength separation between the inner (23∘,30∘) and outer (44∘,50∘) beams, going from Å to Å. A challenge with the 3 MJ HDC design used in this study is that one must overcome radiative losses caused by mixing of the carbon ablator into the hotspot. This requires a higher implosion velocity than necessary for previous designs. Increased velocity provides the necessary hotspot energy and density to overcome bremsstrahlung losses, which was achieved by increasing the ablator thickness and reducing the fuel mass.
Magnetizing the 3 MJ EYC design demonstrates the potential limitations of simple application of external magnetic fields. As seen in Fig. 5 (f), the benefit of magnetization for this design is limited if not detrimental. First of all, we were not able to retrieve the full 25 MJ results from Ref. MacLaren et al., 2024, achieving at best an initial 13.7 MJ yield at T, which steadily degrades with increased magnetic field strength. The primary difference here is that we used laser power multipliers as opposed to opacity multipliers to be consistent with the rest of the simulations in this study. When symmetrized, which is the ideal result, we are able to achieve 30 MJ. We can see the negligible influence of a magnetic field up to 100 T, with slight degradation in the yield until it rises again at 90 T, before falling precipitously again as we increase field strength up to 200 T. The peak in at 90 T suggests that we might obtain some modest benefit of magnetization with some optimization of the design. A noticeable distinction here is that non-trivially decreases with magnetic field strength as seen in Fig. 4.e). The origin of this is not clear yet and is the subject of ongoing investigation.
@@
As yield performance increases, we see diminishing returns from applied magnetic fields. The basic understanding is that the hotspot already gets so hot that it does not heat much more from constrained heat conduction and alpha magnetization. There is a natural peak in the cross section of the DT nuclear reaction at about 50 keV, and we see clearly diminishing returns as one exceeds 20 keV. This trend can be seen in Fig. 17, where we plot the burn fraction, i.e., , versus an analytical modelAtzeni and Meyer-ter-vehn (2004) of burn efficiency.
for the shots in consideration (empty circle), the peak-associated magnetization case (empty star), and the associated symmetrized runs (solid circle and star). Here and is the mass and areal density of the imploded fuel, respectively, as it begins to burn and , where is the sound speed. depends only on and minimizes as 7 g/cm2 for 30 keV. In addition, we have plotted out the burn-efficiency curves for constant values, while for the shots in question we considered the maximum achieved in the simulation. The estimated maximum burn efficiency is 30%, which is being approached by N221204, PSS with low mix, and EYC when magnetized. When designs are already achieving very hot temperatures they will not benefit from magnetization, thus the motivation for designs with lower implosion velocity and higher , such as the PSS design.